Neil eats about 2700 calories per day. On average, he consumes 55.5 grams of fat and 150 grams of protein per day. 

How many of his kcalories are from carbohydrates? 

Please calculate percentages for each macronutrient in order to answer the following: 

What percentage of his diet is made up of carbohydrates? 

What percentage of his diet is made up of protein? 

What percentage of his diet is made up of fat? 

Is Neil within the AMDR for each of the macronutrients? 
Answer #2

Energy in one gram of fat is 9 kcal. Amount of fats consumed is 55.5 grams.

Total energy received in the form of fats is 55.5 * 9 = 499.5 calories and on rounding it off , let’s take it as 500kcal

Energy in one gram of protein is 4 kcal. Amount of protein consumed is 150 grams.

Total energy in the form of proteins is 150*4 = 600 kcal

Total number of calories consumed = 2700

calories in the form of carbohydrates = total energy - ( energy from fats + energy from proteins )

Which is 2700- (600+500) = 1600 kcal.

1600 k calories are from carbohydrates.

———————————————————

Total calories consumed is 2700kcal

Energy from carbohydrates is 1600 kcal.

Percentage of energy in form of carbohydrates is =( energy from carbohydrates * 100 )/ total calories consumed

Which is (1600*100)/2700 = 59.26%

59.26% of his diet is made up of carbohydrates.

Energy from proteins is 600 kcal.

Percentage of energy in form of proteins = (energy from proteins * 100)/ total calories consumed

which is (600*100)/2700 = 22.22 %

22.22% of his diet is in the form of proteins.

Energy from fats is 500 kcal.

Percentage of energy in form of fats = (energy from fats * 100)/ total calories consumed

which is (500*100)/2700 = 18.5 %

18.5% of his diet is in the form of fats.

59.26% of total energy from carbohydrates, 22.22 of total energy from proteins and 18.5% of total energy from fats.

He is little low on fats with 18.5% instead of recommended 20-35%. But it’s a minute deviation and can be acceptable.

By considering the above values, yes Neil is within AMDR for each of the macronutrients.

This is a brief discussion on the given topic. Hope this helps you out. Please feel free to give a like and contact us if you have any further queries. Thank and have a great day.
Answer #1

Total calories consumed per day are 2700kcal.

Amount of fats consumed is 55.5 grams. Energy in one gram of fats is 9kcal.

Total energy consumed in the form of fats is 55.5*9 = 499.5 kcal which is 500 kcal approximately.

Amount of protein is consumed is 150 g. Energy in 1 g of protein is 4Kcal.

Total energy consumed in the form of proteins is 150*4 = 600kcal.

Remaining energy is in the form of carbohydrates which is = total calories consumed - ( energy from fats + energy from proteins)

which is 2700- (500+600) = 1600

1600 of his kcalories are from carbohydrates.

———————————

Total calories consumed in a day are 2700

Energy from carbohydrates is 1600 kcal.

Percentage of total energy from carbohydrates = (energy from carbohydrates*100)/total consumed calories

which is (1600*100)/2700 = 59.26% approximately.

59.26% of his diet is made up of carbohydrates.

Energy from protein is 600kcal

percentage of total energy from proteins = ( energy from proteins * 100) / total calories consumed

which is (600*100)/2700 = 22.22%.

22.22% of his diet is made up of proteins.

Energy from fat is 500 kcal.

Percentage of total energy from fat = ( energy from fats * 100) / total calories consumed

which is (500*100)/2700 = 18.5 %

18.5% of his diet is made up of fats.

With 59.25% of his total energy from carbohydrates, 22.22% from proteins and 18.5% from fats. Neil is little short on fat consumption which is 18.5% of total calorie consumption instead of recommended 20-35%. But it is a very minute deviation and can be acceptable.

By considering above percentages of consumption. Yes, Neil is within AMDR for each of the macronutrients.

This is a brief discussion on the given topic. Hope this helps you out. Please feel free to give a like and contact us if you have any further queries. Thank you and have a great day.